RKY22:

Write an equation in slope-intercept form of the line with the given characteristics through: (-3,1) slop=2

dude:

General slope-intercept form is written as \(y=mx+b\) They give you the slope and a point Points are written as (x,y) First start off by substituting what they gave you

RKY22:

ok

RKY22:

then?

dude:

Once you have that then you can solve for b Do you know how to substitute and solve for it?

RKY22:

no not really im not really in school that much so no i guess not

dude:

Ah Substituting just means switching the variable (or letter for whatever number they give you) In this case, they told you slope = 2 So you would do \(y=2x+b\) Does that make sense?

RKY22:

umm i guess

dude:

Hmm what confuses you?

RKY22:

math

dude:

Haha in specific?

RKY22:

idrk

dude:

Don't think too hard on it You're just switching numbers and letters \(y=mx+b\) y = y value m = slope x = x value b = y-intercept We started out with slope So \(y=mx+b\) becomes \(y=\color{red}2x+b\) Yes?

RKY22:

yes

dude:

Nice Now, we have point \((-3,1)\) Points is the same as saying \((x,y)\) This would mean -3 is our "x value" and 1 is our "y value" Do you know how to substitute them into the equation now?

RKY22:

i think so yeah

dude:

Go ahead and try it (I'll let you know if its good or if it needs a fix)

RKY22:

ok thx